OpenStudy (anonymous):

help!! ill give medal given f(x) = x+7/5 solve for f-1(3)

OpenStudy (anonymous):

f(x) = x+ 7/5 To find inverse, switch places of f(x) and x x= f(x) + 7/5 f(x) = x - (7/5) Now plug in 3 f-1(3) = 3 - (7/5) f-1(3) = ?

OpenStudy (anonymous):

1.6?

OpenStudy (anonymous):

Yes :) 8/5 = 1.6

OpenStudy (anonymous):

the only anwser closest to that on my practice study sheet is 2

OpenStudy (anonymous):

what are the answer choices?

OpenStudy (anonymous):

Wait, for you intital problem is it: \[f(x) = \frac{ x+7 }{ 5}\] or this 2nd one: \[f(x) = x + \frac{ 7 }{ }\]

OpenStudy (anonymous):

x+ (7/5